# Apollo: An Lightweight Multilingual Medical LLM towards Democratizing Medical AI to 6B People

### v2

@@ 4 XMedBench, Multilingual Medical Knowledge Assessment
USMLE MedMCQA MMLUnormal-$\diamondsuit$ MCMLE CMMLUnormal-$\diamondsuit$ FrenchMedMCQA HEAD-QA MMLUnormal-$\diamondsuit$ MMLUnormal-$\diamondsuit$
@@
Results
As shown in Tab. 3, GPT-4 and Qwen-72B rank first in closed source and open source with accuracy rates of 73.37 and 68.74 respectively. The gap between closed source and open source is gradually decreasing. The Apollo series models achieve the best performance of models of the same size. Apollo-7B achieve the same performance as GPT-3.5, Apollo-1.8B achieve the same performance as Mistral-7B, and Apollo-0.5B achieve the same performance as Llama2-7B.
From a language perspective, the models generally does not support Arabic and Hindi well, which further reflects the importance of multilingual medical care. It is worth noting that the two models GPT-4 and Gemma-7B clearly support these two languages better, reflecting the optimization of top open source models for multi-language scenarios. For French, Mistral is the best model of its size. For Spanish, performance is roughly in sync with English. For Chinese, Qwen series models perform better.
At the same time, we also discover some interesting phenomena. BioMistral’s continued training on PubMed-Central based on Mistral does not improve significantly, and the same is true for Meditron. This further confirm that continuing to train extensively on PubMed-Central has little effect and low cost performance, which is mentioned in Sec 3.1.
5 Apollo, the Lite Multilingual Medical LLM
We have two main starting points for training the small model. First, medical equipment usually cannot call network services due to its strict privacy protection settings. For local services, the small model can achieve offline inference on the PC side, which is completely Ensure that the data is localized to help improve the efficiency of medical staff; secondly, the original intention of our article is to explore a reproducible technical solution at an affordable computing cost, and promote the exploration of the field and the raising of new questions. Small models are useful for The training is very friendly for academic researchers who lack sufficient computing power.
Training models in the medical field usually involves continuing pre-training on the corpus. However, some scholars believe that although training on the original corpus gives the model domain knowledge, it greatly damages its ability to prompt question answers (Cheng et al., 2023). We consider exploring ways to rewrite the pre-training corpus into the form of question-and-answer pairs to alleviate this problem (Chen et al., 2023a). At the same time, we use priority sampling methods (Chen et al., 2023a) to achieve a smooth transition between continued pre-training and Instrcution Tuning to ensure the continuity of learning rate and data distribution transformation.
Preliminaries
Our dataset $D$ comprises continuing pre-training data $D_{PT}$ and instruction tuning data $D_{SFT}$. The sampling probability of each data $x\in D$ changes over the course of training. The sampling probability of data $x$ at step $t$ during training was determined using priority sampling, defined as:
$P_{t}(x)=\frac{\pi(x)}{\sum_{y\in D-S_{t}}\pi(y)}$
Settings
@@ 5 Apollo, the Lite Multilingual Medical LLM
USMLE MedMCQA MMLUnormal-$\diamondsuit$ MCMLE CMMLUnormal-$\diamondsuit$ FrenchMedMCQA HEAD-QA MMLUnormal-$\diamondsuit$ MMLUnormal-$\diamondsuit$
@@
Results
As shown in the Tab. 4, under the experimental setting of pre-training first and then SFT, rewriting the pre-training into question and answer has no loss on the overall effect of the model. We also notice that the performance drop after rewriting in English and Hindi, but other languages’ performance improve. After adopting the smooth transition method, we find that except English, other languages’ performance greatly improve. This may be because the data distribution transformation of the previous method is too rigid, resulting in the inability to learn knowledge of long-tail languages (such as Hindi). Using the method of converting to question and answer pairs and making a smooth transition may be able to minimize the knowledge loss of distribution transformation, allowing the model to fully learn the knowledge of long-tail languages and improve the ability of non-mainstream languages. As shown in the Tab. 5, models’ multilingual medical capabilities have been significantly improved after Mix Training.
Model English Chinese French Spanish Arabic Hindi Avg.
Other Models
GPT-3.5 63.81 54.57 68.54 71.48 39.70 39.94 56.34
Meditron-7B 34.83 27.33 24.00 32.81 1.65 18.27 23.15
ProxyTuning for Qwen
Qwen-7B 41.70 49.52 37.69 45.05 28.31 24.89 37.86
Qwen-7B-ProxyTuning 39.83 -${}_{\textbf{-1.87}}$ 51.40 ${}_{\textbf{+1.88}}$ 43.30 ${}_{\textbf{+5.61}}$ 46.97 ${}_{\textbf{+1.52}}$ 29.69 ${}_{\textbf{+1.38}}$ 24.89 ${}_{\textbf{+0.00}}$ 40.79 ${}_{\textbf{+2.93}}$
Table 6: Proxy Tuning for Larger Models
6 The application of the lite Apollo: Proxy Tuning for Larger Models
Preliminaries
Inspired by  Liu et al. (2024a; 2021), we introduce a lightweight model-agnostic decoding method in medical senarios. We leverage the logits from both pre and post fine-tuned small models to indirectly steer the larger base model’s adjustments, thereby eschewing the need for direct parameter fine-tuning. Let $M_{raw}$ denote the smaller pre-trained model, and $M_{tuned}$ denote its fine-tuned counterpart. We compute the logit offset as ”proxy” for each token, corresponding to the anti-expert and expert roles as delineated in  Liu et al. (2021). This offset is then applied to the base model $M_{base}$ to synchronize the predictive distributions of the smaller and larger models. The modified probability distribution is given by:
$p^{\prime}_{\mathcal{M}_{base}}(X_{t}\mid x_{1,...,t-1})=\text{softmax}\left[l_{M_{base}}+\Delta l_{M}\right]\propto p_{\mathcal{M}_{base}}(X_{t}\mid x_{1,...,t-1})\left(\frac{p_{\mathcal{M}_{tuned}}(X_{t}\mid x_{1,...,t-1})}{p_{\mathcal{M}_{raw}}(X_{t}\mid x_{1,...,t-1})}\right)$
Settings
$M_{base}$ is designated as the subject of investigation for Qwen-7B. Apollo-1.8B and Qwen-1.8B are appointed as the $M_{tuned}$ and $M_{raw}$ respectively.
Results
As shown in the Tab. 6, the overall effect of the model improves a lot without changing the parameters after Proxy Tuning. From language perspective, except English, all other languages increase, and French has the most obvious increase. Excitingly, for French and Spanish, the model after Proxy Tuning performs better than both $M_{tuned}$ and $M_{base}$, indicating that new accurate knowledge is generated after Proxy Tuning. We also notice a decline in English proficiency. This may be because there is a gap between the distribution of difference and the probability itself, which leads to over-strengthening of the second option and requires further exploration and optimization.
7 Related Work
[…]
Notable early developments in healthcare-focused LLMs include GatorTron (Yang et al., 2022), which explores the use of unstructured Electronic Health Records (EHRs) in clinical settings, and Codex-Med (Liévin et al., 2022), which assesses the efficacy of GPT-3.5 models in healthcare. Galactica (Taylor et al., 2022) targets the challenge of scientific information overload by storing, combining, and reasoning about scientific knowledge, including in healthcare.
Privacy and confidentiality concerns have prompted innovative solutions like DeID-GPT (Liu et al., 2023), a de-identification framework using GPT-4, and MedAlpaca (Han et al., 2023), which offers an open-source, on-site implementation approach. ChatDoctor (Yunxiang et al., 2023), Meditron (Chen et al., 2023b) and PMC-LLaMA (Wu et al., 2023) address limitations in medical knowledge accuracy of existing LLMs, tuning base models on millions of biomedical papers.
GatorTronGPT (Peng et al., 2023), a clinical LLM with a GPT-3 architecture, and MedAGI (Zhou et al., 2023), a model aiming to unify domain-specific medical LLMs, illustrate the trend towards more specialized models. Finally, Med-Flamingo (Moor et al., 2023) represents an advance in vision-language models, handling multimodal data that includes both images and text.
8 Conclusion
In order to serve more people and larger community, we carefully collect and organize a high-quality medical corpus covering most populous languages in the world, open sourcing multi-language Dataset ApolloCorpora and evaluation set XMedBench. Based on these, we explore suitable methods for multilingual training and interrelationships between languages in the medical field, and finally obtains a series of models named after Apollo, with SOTA performance from 0.5B to 7B. At the same time, the Proxy Tuning method is also used to improve larger models’ multilingual medical capabilities without changing the parameters.
We propose the question of exploring the relationship between languages in the medical field. While open-sourcing a series of small models with comparable performance, we expand their application scenarios and keep the entire process open-source and reproducible. We provide a foundation for researchers with limited resources to explore questions such as datasets sampling optimization, Proxy Tuning optimization, and different language models combination optimization in the multilingual medical field.
Acknowledgements
This work is supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), and Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608).
We recognize the existence of a concurrent work by Qiu et al. (2024) (Qiu et al., 2024), which became publicly available prior to our work. It is important to note, however, that our research was conducted before the release of Qiu et al.’s work, with the interval before our publication being devoted to thorough review and refinement of our manuscript.
Appendix A Prompts
A.1 Prompt for Generating Doctor-Patient Dialogues
Figure 5: Prompt Template for Generating Doctor-Patient Dialogues
A.2 Prompts for Generating QA Pairs from Texts
Appendix B Models
Qwen

### v4

@@ 3.2 Apollo, the Lite Multilingual Medical LLM
[…]
@@
Settings We set the priority $\pi(x)=16$ for $x\in D_{PT}$, and $\pi(x)=2$ for $x\in D_{SFT}$. In order to achieve the purpose of smooth transition of sampling ratio. The overall training sequence of pre-training corpus first, and then instruction Tuning corpus is maintained, but the transition can be smoothed. The Batch size of model training is set to 256, the learning rate is set to 1e-5, and the warm up rate of Cosine scheduler is set to 0.03. The pre-training corpus is trained for one epoch, the instruction data is trained for two epochs.
4 Evaluation
4.1 XMedBench: Multilingual Medical Knowledge Evaluation
We focus on assessing multilingual medical knowledge, select multiple-choice questions as tasks, and collect common data sets with local medical characteristics, see details in App. B.
Construction of XMedBench For English, we use the MedQA-USMLE (Zhang et al., 2018), MedMCQA (Pal et al., 2022), and medical-related parts of MMLU (Hendrycks et al., 2020); for Chinese, we used the of MedQA-MCMLE (Zhang et al., 2018) and medical-related parts CMMLU (Li et al., 2023a); for Spanish, we used HEAD-QA (Vilares and Gómez-Rodríguez, 2019); for French, we used FrenMedMCQA (Labrak et al., 2023a); For Arabic and Hindi, which lack local evaluation, we compromised and followed Llama3’s multilingual evaluation method (et al., 2024), using Google Translate and inviting practicing physicians to proofread and finally get the translated version of MMLU. Specifically, we follow Med-PaLM2 (Singhal et al., 2023) and select six subcategories in MMLU: Clinical knowledge, Medical genetics, Anatomy, Professional medicine, College biology, and College medicine. For MedQA, we choose the 4-options version. For CMMLU, we select seven subdirectories: Anatomy, Clinical knowledge, College medicine, Genetics, Nutrition, Traditional chinese medicine, and Virology.
Settings We adopt 3-shot evaluation and use regular matching to extract options. The specific evaluation prompts are shown in Fig. 3. For the generation strategy, we do not perform sampling and set the maximum and minimum number of generated tokens to 128 and 2. For model loading, except for the 0.5B size model which uses full precision loading, we uniformly use half precision loading. Please see the App. B.2 for details of Models.
@@ 4.1 XMedBench: Multilingual Medical Knowledge Evaluation
4.2 Benchmarking results
As shown in Tab. 3, GPT-4 and Qwen-72B rank first in closed source and open source with accuracy rates of 73.37 and 68.74 respectively. The gap between closed source and open source is decreasing. The Apollo series models achieve the best performance of models of the same size. Apollo-7B achieve comparable performance as GPT-3.5, Apollo-1.8B achieve comparable performance as Mistral-7B, and Apollo-0.5B achieve comparable performance as Llama2-7B.
From a language perspective, all models scored worse on Arabic and Hindi compared to English, which further demonstrates the medical community’s neglect of these two languages. Note that GPT-4 support these languages better, reflecting OpenAI’s emphasis on multi-language scenarios. Mistral is better adapted to French, and the Qwen and Yi models have better support for Chinese.
4.3 More Analysis
As shown in the Tab. 4, under the experimental setting of pre-training first and then SFT, rewriting the pre-training into question and answer has no loss on the overall effect of the model. We also notice that the performance drop after rewriting in English and Hindi, but other languages’ performance improve. After adopting the smooth transition method, we find that except English, other languages’ performance greatly improve. This may be because the data distribution transformation of the previous method is too rigid, resulting in the inability to learn knowledge of long-tail languages (such as Hindi). Using the method of converting to question and answer pairs and making a smooth transition may be able to minimize the knowledge loss of distribution transformation, allowing the model to fully learn the knowledge of long-tail languages and improve the ability of non-mainstream languages.
As shown in the Tab. 6, models’ multilingual medical capabilities have been significantly improved after Mix Training. For different model sizes, although the improvement effect gradually decreases as the model parameters increase, the model performance still continues to increase, indicating promising prospects for scaling up training. Impressively, Apollo-0.5B achieves considerable performance with few parameters. Given its potential for real-time inference on a wide range of hardware, we believe it can democratize advances in medical AI to the broader community.
Model English Chinese French Spanish Arabic Hindi Avg.
Other Models
GPT-3.5 63.81 54.57 68.54 71.48 39.70 39.94 56.34
Meditron-7B 34.83 27.33 24.00 32.81 1.65 18.27 23.15
Proxy-Tuning for Qwen
Qwen-7B 41.70 49.52 37.69 45.05 28.31 24.89 37.86
Qwen-7B-Proxy-Tuning 39.83 -${}_{\textbf{-1.87}}$ 51.40 ${}_{\textbf{+1.88}}$ 43.30 ${}_{\textbf{+5.61}}$ 46.97 ${}_{\textbf{+1.52}}$ 29.69 ${}_{\textbf{+1.38}}$ 24.89 ${}_{\textbf{+0.00}}$ 40.79 ${}_{\textbf{+2.93}}$
Table 5: Proxy-Tuning for Larger Models
5 The Application of the Lite Apollo: Proxy-Tuning for Larger Models
Preliminaries Inspired by  Liu et al. (2024a, 2021), we introduce a lightweight model-agnostic decoding method in medical senarios. We leverage the logits from both pre and post fine-tuned small models to indirectly steer the larger base model’s adjustments, thereby eschewing the need for direct parameter fine-tuning. Let $M_{raw}$ denote the smaller pre-trained model, and $M_{tuned}$ denote its fine-tuned counterpart. We compute the logit offset as "proxy" for each token, corresponding to the anti-expert and expert roles as delineated in  Liu et al. (2021). This offset is then applied to the base model $M_{base}$ to synchronize the predictive distributions of the smaller and larger models. The modified probability distribution is given by:
$p^{\prime}_{\mathcal{M}_{base}}(X_{t}\mid x_{1,...,t-1})=\text{softmax}\left[l_{M_{base}}+\Delta l_{M}\right]$
$\propto p_{\mathcal{M}_{base}}(X_{t}\mid x_{1,...,t-1})\left(\frac{p_{\mathcal{M}_{tuned}}(X_{t}\mid x_{1,...,t-1})}{p_{\mathcal{M}_{raw}}(X_{t}\mid x_{1,...,t-1})}\right)$
Settings $M_{base}$ is designated as the subject of investigation for Qwen-7B. Apollo-1.8B and Qwen-1.8B are appointed as the $M_{tuned}$ and $M_{raw}$ respectively.
Results As shown in the Tab. 5, the overall effect of the model improves a lot without changing the parameters after proxy-tuning. From language perspective, except English, all other languages increase, and French has the most obvious increase. Excitingly, for French and Spanish, the model after proxy-tuning performs better than both $M_{tuned}$ and $M_{base}$, indicating that new accurate knowledge is generated after proxy-tuning. We also notice a decline in English proficiency. This may be because there is a gap between the distribution of difference and the probability itself, which leads to over-strengthening of the second option and requires further exploration and optimization.
6 Related Work
[…]
From a language perspective, many excellent works have appeared in the Chinese and English medical fields respectively. For Chinese, HuatuoGPT (Chen et al., 2023a) and BenTsao (Wang et al., 2023a) achieved good results by training on Chinese wikis, papers and medical consultation data. For English, Meditron (Chen et al., 2023b) and PMC-LLaMA (Wu et al., 2023) address limitations in medical knowledge accuracy of existing LLMs, tuning base models on millions of biomedical papers. For other languages, to the best of our knowledge, corresponding Medical LLMs have not yet emerged.
7 Conclusion
In order to serve more people and larger community, we carefully collect and organize a high-quality medical corpus covering most populous languages in the world, open sourcing multi-language Dataset ApolloCorpora and evaluation set XMedBench. Based on these, we explore suitable methods for multilingual training and interrelationships between languages in the medical field, and finally obtains a series of models named Apollo, with SOTA performance from 0.5B to 7B. Meanwhile, proxy-tuning is used to improve large foundation in terms of multilingual medical capabilities without changing the parameters. We offer a foundation for global researchers, specially those with limited resources, to investigate medical LLMs.
8 Acknowledgement
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing.
Appendix A Details of ApolloCorpora, Multilingual Medical Dataset
A.1 Dataset Taxonomy and Collection of ApolloCorpora
Books For English books, we use medical dictionary  [^3] https://www.nlm.nih.gov/research/umls/new_users/online_learning/LEX_001.html to filter the books in the Pile Dataset (Gao et al., 2020) and select books with medical words accounting for more than 4%, and finally obtain 2312 medical-related books. For Chinese books, we follow MedQA (Jin et al., 2020) to collect medical textbooks included in the five-year and eight-year medical student training programs in mainland China, and finally obtain 90 books.
Papers For English papers, we sample the public data in PubMed and obtain 878,241 medical abstracts. For Chinese papers, we also screen a total of 177,261 abstracts of papers published by the Chinese Medical Association  [^4] https://www.yiigle.com/index. For French papers, we use the MORFITT (Labrak et al., 2023b) dataset and the scientific article portion of the CLEAR (Grabar and Cardon, 2018). For the Spanish paper, we use paper abstracts open sourced by the Mesinesp (Gasco et al., 2021).
Figure 5: Prompt Template for Generating Doctor-Patient Dialogues
Encyclopedias For the English Encyclopedia, we also use the English Medical Dictionary to filter out 36107 medical-related wiki pages from dataset [^5] https://huggingface.co/datasets/wikipedia. For the French encyclopedia, we select the encyclopedia articles part of the CLEAR (Grabar and Cardon, 2018). For the Hindi encyclopedia, we choose the HHD corpus (Jain and Arora, 2018), which crawls descriptions of people, diseases, medical consumer products, and symptoms from Indian websites.
Doctor-Patient Dialogues For Chinese, we directly use the HuatuoGPT dataset (Zhang et al., 2023) and the simplified data set in Huatuo_26M (Li et al., 2023b). For English, we construct a multi-turn conversation data set based on PMC-Patients (Zhao et al., 2022) using ChatGPT, Prompt is shown in the Fig. 5. For Arabic, we extract high-quality questions and answers with both question and answer lengths greater than 128 from the largest Arabic healthcare question and answer dataset MAQA (Abdelhay and Mohammed, 2022).
Exams For the Chinese exam, we collect training sets of CMB (Wang et al., 2023b), CMExam (Liu et al., 2024b), and MedQA (Zhang et al., 2018). For the English exam, we collect the training sets of MedQA, Medmcqa (Pal et al., 2022) and Pubmedqa (Jin et al., 2019). For the Spanish and French exam, we select the training set of HEAD-QA (Vilares and Gómez-Rodríguez, 2019) and Frenchmcqa (Labrak et al., 2023a) separately.
Guidelines For English Guidelines, we select data from three sub-items of NICE [^6] https://www.nice.org.uk/guidance, PubMed and SPOR [^7] https://sporevidencealliance.ca/key-activities/cpg-asset-map/cpg-database/ in the clinical guidelines introduced by Meditron (Chen et al., 2023b).
General Instruction Tuning We use the translation (FreedomIntelligence, 2023) and original data of Sharegpt [^8] https://github.com/lm-sys/FastChat and Alpaca (Taori et al., 2023). For Chinese, we additionally make use of data (FreedomIntelligence, ) generated by GPT-4 based on WizardLM Method (Xu et al., 2023). For English, in addition to adding the WizardLM Dataset, we also add belebele (Bandarkar et al., 2023) to enhance multi-language reading comprehension capabilities, ai2_arc (Clark et al., 2018) to enhance abstract reasoning capabilities, Capybara (Daniele and Suphavadeeprasit, 2023) to enhance instruction following capabilities.
Web For Chinese, we use the medical dictionary (Han et al., 2016) to filter out medical-related articles from the Wudao Dataset (Xue et al., 2022). For English, we use the English Medical Vocabulary [^9] https://www.nlm.nih.gov/research/umls/new_users/online_learning/LEX_001.html to filter out medical related articles in C4 Dataset (Raffel et al., 2019). For Spanish, we sampled 10% of CoWeSe Dataset (Carrino et al., 2021). Math For mathematical abilities, we choose MathInstruct (Yue et al., 2023), a composite dataset containing various mathematics-related tasks and problem formats. Code We choose Python-Alpaca (Vezora, 2023) and Leetcode-ZH-11k (krisfu, 2023) respectively to strengthen the ability to solve coding tasks in Chinese and English.
Examples Chinese: Representatives of exogenous wind-cold and high fever include: (A) Yinqiao Powder, (B) Qingwen Baidu Decoction, (C) Dachengqi Decoctio. (D) Xiaobuhu Decoction Arabic: Patient: “Peace be upon you, I feel very intense friction in my body and head, redness in the skin or its color changing to pink, and I am almost cut off in my body. This all happens when I start making an effort, even if it is the smallest of things, even if my voice rises or I get angry. Thank you, and may God’s peace, mercy, and blessings be upon you.” Doctor: “Often this is sensitivity due to exertion and sweating, and they are signs of a disease called fever of the nile, or a rash or allergy to sweating that appears in the summer season in particular. It is preferable to use a cold shower and apply the pimples with a weakening ointment. If the area is small, apply Calcipotriene ointment to it, take an anti-histamine medication such as Polaramine." Hindi: Disease: Epilepsy. Tulsi reduces many diseases like a panacea. A large amount of anti-oxidants are found in Tulsi which cures free radicals in the brain. In case of any type of brain disease, if taken daily If 20 basil leaves are chewed and eaten, it is very effective. Brahmi leaves: Brahmi leaves grow around our homes especially where there is soil. It is round and curved in shape. It is taken daily. Consuming it chewed on an empty stomach not only strengthens the memory but also reduces epileptic seizures. Spanish: Due to its high mercury content, the Spanish Agency for Food Safety and Nutrition recommends pregnant women not consume: A. Ripe soft cheeses, such as Brie or Camembert, B. Pâté or foie-gras, C. Calmette and Guérin bile vaccine., D. Raw sausages French: What vaccinations are required in France? A.Measles-Mumps-Rubella, B.Human papillomavirus, C.Diphtheria-Tetanus-Poliomyelitis, D.Whooping cough, E.Calmette and Guérin bile vaccine.
A.2 Details for Data Rewriting of ApolloCorpora
We want to explore whether rewriting the original pre-training corpus into QA pairs in the context of continuing training can help increase its medical capabilities without destroying the original model’s capabilities. We use ChatGPT101010gpt-3.5-turbo-16k-0613 to generate questions and answers for a certain paragraph. For paragraph interception, we divide it according to the basic semantic units in the data set, such as sections in books and guides, paragraphs in website data, single wiki entry and abstracts of papers. For basic semantic units that are too long, we comprehensively consider the knowledge expression density of the language and subdivide different languages into blocks of different lengths to ensure that the semantic information covered by a single paragraph does not exceed the amount of information that can be included in a question and answer pair. For Spanish, French, English and Hindi we use 2048, for Chinese we use 256 and for Arabic we use 128. Prompts for generating QA pairs are detailed in the Fig. 5 and Fig. 4.
A.3 Localized features of ApolloCorpora
In terms of symptom diagnosis, local languages retain the terminology of traditional medicine, and due to different geographical environments and living habits, the possibility that a certain symptom corresponds to different diseases is also different: for Chinese, a disease has two aspects: "bìng" and "zhèng". The former is often translated as "disease entity". The latter, and more important one, is usually translated as "pattern". For example, the disease entity of a common cold might present with a pattern of wind-cold in one person, and with the pattern of wind-heat in another [^11] https://en.wikipedia.org/wiki/Traditional_Chinese_medicine#Six_Excesses.
In terms of communication terms, some languages will have religious-related idioms at the beginning and end to improve the communication experience, such as Arabic.
@@ A.3 Localized features of ApolloCorpora
[…]
@@
Appendix B Details of XMedBench
B.1 Construction of XMedBench
For English, we use the MedQA-USMLE (Zhang et al., 2018), MedMCQA (Pal et al., 2022), and medical-related parts of MMLU (Hendrycks et al., 2020); for Chinese, we used the of MedQA-MCMLE (Zhang et al., 2018) and medical-related parts CMMLU (Li et al., 2023a); for Spanish, we used HEAD-QA (Vilares and Gómez-Rodríguez, 2019); for French, we used FrenMedMCQA (Labrak et al., 2023a); For Arabic and Hindi, which lack local assessments, we make a compromise by applying translated versions of MMLU [^12] Hindi: https://huggingface.co/datasets/FreedomIntelligence/MMLU_Hindi; Arabic: https://huggingface.co/datasets/FreedomIntelligence/MMLU_Arabic. Specifically, we follow Med-PaLM2 (Singhal et al., 2023) and select six subcategories in MMLU: Clinical knowledge, Medical genetics, Anatomy, Professional medicine, College biology, and College medicine. For MedQA, we choose the 4-options version. For CMMLU, we select seven subdirectories: Anatomy, Clinical knowledge, College medicine, Genetics, Nutrition, Traditional chinese medicine, and Virology.
B.2 Models for XMedBench